Using thye following information: H0:=375; Ha: 375; n=100; x = 390; 2=4,900 and a =0.05. is the z-test statistic value equal to 2.14?


1
Expert's answer
2022-05-19T11:12:36-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=375"

"H_1:\mu\not=375"

Based on the information provided, "n=100, \bar{x}=390, \sigma^2=4900," the significance level is "\alpha = 0.05."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{390-375}{\sqrt{4900}/\sqrt{100}}=\dfrac{15}{7}\approx2.14"